# supplement.pdf

### v1

@@ Abstract
To provide insight into the basic properties of emerging structures when bacteria or other microorganisms conquer surfaces, it is crucial to analyze their growth behavior during the formation of thin films. In this regard, many theoretical studies focus on the behavior of elongating straight objects. They repel each other through volume exclusion and divide into two halves when reaching a certain threshold length. However, in reality, hardly any object of a certain elongation is perfectly straight. Therefore, we here study the consequences of curvature on the growth of colonies and thin active films. A given amount of curvature is prescribed to each growing individual. Particularly, we analyze how this individual curvature affects the size of orientationally ordered domains in the colony and find a significant decrease. Instead, strings of stacked curved cells emerge that show branched structures. Furthermore, we identify a significant spatio-orientational coupling that is not observed in colonies of straight cells. Our results are important for a fundamental understanding of the interaction and spreading of microorganisms on surfaces, with implications for medical applications and bioengineering.
Figure 1: Illustration of our curved model microorganisms. (A) Each microorganism $i$ is represented by a curved spherocylinder of thickness $d_{0}$ following an arc of radius $R$ and segment angle $\phi_{i}$. We characterize its direction by the unit vector $\hat{\bm{e}}_{i}$ pointing in the direction from the center-of-mass position $\bm{r}_{i}$ radially outward, normal to the arc. The angle $\theta_{i}$ measures the orientation of $\hat{\bm{e}}_{i}$ in the two-dimensional plane. $\bm{r}_{i}^{\mathrm{o}}$ denotes the center-of-arc position for cell $i$. (B) An illustration of calculating the mutual forces of steric interaction between two curved microorganisms. For any two cells $i$ and $j$ in contact, these forces are determined from the overlap of the nearest inner tangent circles of the curved spherocylinders, with center-to-center vector $\bm{d}_{ij}$, using the Hertzian elastic model. (C–E) Illustrative examples of elongated cells for curvatures of (C) $0.011/R=$ 0.01, (D) $101/R=$ 10, and (E) $0.331/R=$ 0.33. (F) A typical time evolution of the total number of cells $N(t)$ of curvature $0.101/R=$ 0.10 and threshold division length $L_{\mathrm{d}}=5$. The star symbols indicate configurations of $N(t)=2$, $115$, and $1031$, as depicted in panel (G). Color represents the orientation angle $\theta_{i}$ of the cells.
@@
To this end, we introduce a theoretical model to study the growth of the films. It is based on discrete growing mesoscopic curved objects that divide into two halves when their arc length reaches a certain threshold. We illustrate the effects of curvature of the individual microorganisms on the structure of the resulting films. Particular attention is paid to the average size of ordered domains and how it depends on the curvature of the individuals. Moreover, the influence of the threshold length of cell division is investigated. We also discuss the statistical properties of the orientation of curved cells and report a spatio-orientaional coupling that we do not observe for straight objects. In addition, we briefly address different structures emerging for growing curved objects.
Results and Discussion
Model – growing and dividing curved spherocylinders
@@
Figure 2: Effects of curvature $1/R$ of the growing individual cells and their threshold division length $L_{\mathrm{d}}$ on the resulting overall structure of the colony. From left to right, curvature increases as $0.011/R=$ 0.01, $0.10$, and $0.33$. From top to bottom, the threshold division length decreases as $L_{\mathrm{d}}=5$, $4$, and $3$. Increasing curvature and decreasing threshold division length lead to finer, less ordered structures and decreasing average domain size. Color represents the angle of orientation $\theta_{i}\in[0,2\pi)$ of each cell $i$, see Fig. 1A, as quantified by the color bar on the right-hand side. The mean growth rate is set to $g=5\times 10^{-5}$.
@@
Figure 3: (A) Average domain area $\expectationvalue{A}$ as a function of cellular curvature $1/R$ for threshold division lengths $L_{\mathrm{d}}=5$, $4$, and $3$. The average is taken over all domains located within the inner circle of half the colony radius $r_{\mathrm{colony}}$. (B) Dependence of the average domain area for different cellular curvatures $0.011/R=$ 0.01,$0.10$, and $0.33$ on the distance $r$ from the center of the colony. The threshold division length is set to $L_{\mathrm{d}}=5$. We calculate the average $\expectationvalue{A}_{\Delta r}$ for each value of $r$ from all domains located within an annular ring of width $\Delta r=5$. Typically, $r_{\mathrm{colony}}\sim 100$ at times of evaluation. (C) Probability distribution $P(A)$ of the domain area $A$ within colonies of cellular curvature $0.011/R=$ 0.01, $0.10$, and $0.33$. Dashed lines represent exponential behavior $P(A)\sim\exp[-A/A_{*}]$ within the indicated regimes. From corresponding fits, we obtain values of characteristic domain areas $A_{*}=145.4$, $35.2$, and $5.7$, respectively, which decrease with increasing curvature. We set the threshold division length to $L_{\mathrm{d}}=5$. The average is taken over 30 samples.
@@
As outlined above, straight-growing bacteria are frequently represented by growing spherocylinders that divide into two halves after reaching a certain threshold length [8]. To now consider curved microorganisms, we add curvature to these spherocylinders. For this purpose, each curved cell is described as an arc (part of a circle) of radius $R$ and thickness $d_{0}$, see Fig. 1A. The length of the $i$-th cell is given by the arc length $L_{i}=R\phi_{i}$, where $\phi_{i}$ is the opening angle of the segment of the disk containing the microorganism at its bounding arc. Figures 1C, D, and E show illustrations of curved cells of curvature $1/R=0.01,0.10,$ and $0.33$.
@@
During the entire process of growth, we need to evaluate the mutual steric interactions resulting from volume exclusion between the cells and their consequences. To this end, at each time step, we find, for any two cells $i$ and $j$ in contact, the two inner tangent circles to the curved spherocylinders that are nearest to each other [14], see Fig. 1B. The center-to-center vector between these two circles is termed $\bm{d}_{ij}$. We assume the force of mutual interaction $\bm{F}_{ij}$ between the two circles to be described by the Hertzian elastic model [17]. The force $\bm{F}_{ij}$ is then mapped to the interaction between the $i$-th and $j$-th cells.
@@
Here, $\bm{r}_{i}(t)$ denotes the position of the center of mass of the $i$-th cell, and $\theta_{i}(t)$ quantifies the orientation of the $i$-th cell, both at time $t$. The direction of the orientation of the $i$-th cell $\hat{\bm{e}}_{i}=(\cos\theta_{i},\sin\theta_{i})$ is set from the center of the arc to the center-of-mass position of the cell, normal to the arc length, see Fig. 1A. $\theta_{i}$ takes values in the range $[0,2\pi)$ for curved cells, as curvature provides a sense. $\zeta$ sets the strength of the friction with the substrate per arc length of the cell and is supposed to be identical for each microorganism. Moreover, $\bm{\ell}_{ij}(t)$ is the vector from the center-of-mass position to the point of the cell where the force acts. $\hat{\bm{e}}_{z}$ is the unit vector pointing into the $z$-direction, normal to the plane.
@@
Figure 4: (A) Schematic illustration of evaluating the spatio-orientational coupling. We determine for each cell $i$ the angle $\gamma_{i}$ for the radial direction running from the center of mass of the colony through the center of mass of the cell $\bm{r}_{i}$. This angle is compared to the individual angular orientation $\theta_{i}$ for each cell. Histograms of $\theta_{i}-\gamma_{i}$ for cellular curvatures (B) $0.011/R=$ 0.01, (C) $0.101/R=$ 0.10, and (D) $0.331/R=$ 0.33 are depicted for the threshold division length $L_{\mathrm{d}}=5$. In (C) and (D), the dashed lines are obtained by fitting a Gaussian distribution $a\exp[(\theta-\gamma)^{2}/(2\sigma^{2})]+b$ to the histograms. Resulting (E) standard deviations $\sigma$ and (F) amplitudes $a$ of the Gaussian fits are depicted as a function of cellular curvature $1/R$ for the different threshold division lengths $L_{\mathrm{d}}=5$, $4$, and $3$.
@@
On this basis, we now explore the fundamental effects of growing and dividing curved microorganisms on the resulting film structures. Videos of the simulated growth process starting from a single, randomly oriented cell under the aforementioned conditions are available in the Supporting Information. Figure 1F shows a typical resulting time evolution of the total number of cells $N(t)$ in a growing colony, while Fig. 1G provides three intermediate snapshots. Final example configurations upon variation of the basic parameters, that is, curvature $1/R$ (inverse radius, see Fig. 1A) and threshold division length $L_{\mathrm{d}}$, are depicted in Fig. 2. All quantities described in the following are analyzed in the final stages of the simulations.
@@
In Fig. 2, color represents the angle of the orientational vector $\theta_{i}\in[0,2\pi)$ of the $i$-th cell, see Fig. 1A. The columns from left to right show increasing cellular curvature $1/R$. In the left-most column, for the smallest curvature $0.011/R=$ 0.01, the individual cells are basically straight, see Fig. 1C, and we qualitatively expect corresponding behavior. As already reported in Ref. [8] for straight cells, one does not find global orientational order. Instead, locally aligned domains form a kind of “mosaic” structure. These mesoscopic domains decrease in size as the division length $L_{\mathrm{d}}$ decreases. We recover this trend from top to bottom when we reduce this length.
@@
Our central observation and message is that increasing curvature $1/R$ and decreasing threshold division length $L_{\mathrm{d}}$ generally and substantially reduce the average size of the mesoscopic domains and thus the extent of local alignment. This effect becomes apparent from Fig. 2 when turning from left to right (curvature $1/R$ increasing as $0.01$, $0.10$, $0.33$) and from top to bottom (threshold division length $L_{\mathrm{d}}$ decreasing as $5$, $4$, $3$).
@@
In the following, we analyze these effects more quantitatively. To this end, we determine the average area $\langle A\rangle$ of the mesoscopic domains at any given state of the colony. To determine the sizes of the domains of orientational order, we consider two neighboring cells $i$ and $j$ as belonging to the same mesoscopic domain, if they contact each other and show approximately identical orientation. The first criterion is satisfied if the force in (4) becomes nonzero. The second criterion applies if their orientation angles fulfill |$\absolutevalue{\theta_{i}-\theta_{j}}<0.1$. The area $A$ of any resulting mesoscopic domain is calculated as the areal sum over all cells within that domain. We evaluate $\langle A\rangle$ as a function of cellular curvature. Ensemble averages are taken over $10$ different simulations. Figure 3A shows the average domain area $\expectationvalue{A}$ as a function of the cellular curvature $1/R$ for the different threshold division lengths $L_{\mathrm{d}}=5,4,$ and $3$. To exclude surface effects, $\expectationvalue{A}$ is calculated in each case for the interior circular film area of half the colony radius $r_{\mathrm{colony}}/2$. Here, the radius of the entire colony is defined via $r_{\mathrm{colony}}(t)=\max_{i=1,...,N}|\bm{r}_{i}(t)|$. The origin is set by the position of the initial cell. In our case, $r_{\mathrm{colony}}\approx 100$, but the precise value slightly depends on the curvature and division length (see Section 1 of the Supporting Information). From Fig. 3A, we find that the average domain area $\expectationvalue{A}$ is almost constant when increasing the curvature of the cells $1/R$ at small values, that is, for basically straight microorganisms. Yet, from a certain value of $1/R$ onwards, $\expectationvalue{A}$ significantly decreases. This effect is most pronounced for $L_{\mathrm{d}}=5$, while for $L_{\mathrm{d}}=4$ and $3$ it is more moderate. Moreover, the values at large curvature $1/R$ approach each other, and all tend towards the same value. Apparently, in this regime, the average domain size $\expectationvalue{A}$ is rather controlled by curvature than the length of cell division $L_{\mathrm{d}}$.
@@
Next, we address in Fig. 3B how the area of the domains depends on their location within the colony, together with the effect of cellular curvature on this dependence. We slice the entire colony into concentric, annular rings of width $\Delta r$ for quantification. Their location is given by their distance $r$ from the center of the colony. Figure 3B shows the averaged domain area within each ring, $\expectationvalue{A}_{\Delta r}$, as a function of $r$. The ensemble average is taken over 30 different simulations. For basically straight cells of curvature $0.011/R=$ 0.01, we observe a decreasing trend of $\expectationvalue{A}_{\Delta r}$ towards the outer rim of the colony, in line with previous reports in Ref. [8]. Instead, for the increasingly curved cases of $0.101/R=$ 0.10 and $0.33$, $\expectationvalue{A}_{\Delta r}$ remains rather constant towards the outer rim. In line with our results above, we infer that increased curvature alone already implies a significant reduction in domain area and, thus, an increase in mesoscopic disorder. As remarked in Ref. [8] for straight cells, reducing the division length $L_{\mathrm{d}}$ has similar effects. It likewise smoothens the distribution of domain area from the center towards the outer rim of the colony.
@@
Additionally, we evaluate the probability distribution $P(A)$ for the observed mesoscopic domain areas $A$ in a given colony. The distribution shows exponential behavior for straight cells $P(A)\sim\exp[-A/A_{*}]$, where $A_{*}$ represents a characteristic domain area. $A_{*}$ decreases together with the division length $L_{\mathrm{d}}$ [8, 36]. We infer from Fig. 3C that the exponential dependence can still be observed with increasing cellular curvature $1/R$. However, increasing curvature reduces the characteristic area $A_{*}$, in line with our qualitative observation in Fig. 2.
@@
Figure 5: Center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ of the microorganisms in the film, see Fig. 1(A), together with color coding of the individual orientations for increasing cellular curvature (A) $0.011/R=$ 0.01, (B) $0.101/R=$ 0.10, and (C) $0.331/R=$ 0.33. The threshold division length is set to $L_{\mathrm{d}}=5$ so that the illustrations reconsider the state of the colonies depicted in the top row of Fig. 2. For the almost straight cells in (A), the radii of the arcs are correspondingly large, therefore several center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ are located outside the box. As curvature $1/R$ increases from left to right, the mutual arrangements of $\bm{r}_{i}^{\mathrm{o}}$ transition from (A) rather scattered, gas-like to (C) organized, branched filamentous structures. The latter indicates the stacked organization of curved cells in these colonies.
Thus, as demonstrated above, the curvature of individual cells can have an effect qualitatively similar to reduced division lengths. At least, this concerns the average mesoscopic domain size in the colonies. Additional, new structural consequences of cellular curvature emerge that are not present in colonies of straight cells. A specific example is the coupling of cellular orientation to the position within the colony. Through these measures, we can qualitatively and quantitatively distinguish the appearance of colonies of straight cells of low division length from films of curved cells of longer division length, which by eye appear relatively similar; see the bottom left and top right in Fig. 2.
@@
To illustrate this trend, we evaluate their spatio-orientational coupling. We determine for each cell the angle $\gamma_{i}$ of the line running from the center of the colony to the center of mass of the cell. That is, $\gamma_{i}$ measures the radial direction of the colony at the position of the considered cell. We compare this angle with the angular orientation $\theta_{i}$ of each cell; see Fig. 4A. Proceeding along these lines for all cells, we calculate histograms as a function of $\theta_{i}-\gamma_{i}$. The trend with increasing curvature $0.011/R=$ 0.01, $0.10$, to $0.3$ is illustrated in Fig. 4BD. The threshold division length here is set to $L_{\mathrm{d}}=5$. Ensemble averages over ten colonies are performed for each curvature.

### v2

@@ Abstract
To provide insight into the basic properties of emerging structures when bacteria or other microorganisms conquer surfaces, it is crucial to analyze their growth behavior during the formation of thin films. In this regard, many theoretical studies focus on the behavior of elongating straight objects. They repel each other through volume exclusion and divide into two halves when reaching a certain threshold length. However, in reality, hardly any object of a certain elongation is perfectly straight. Therefore, we here study the consequences of the curvature of individuals on the growth of colonies and thin active films. This individual curvature, so far hardly considered, turns out to qualitatively affect the overall growth behavior of the colony. Particularly, strings of stacked curved cells emerge that show branched structures, while the size of orientationally ordered domains in the colony is significantly decreased. Furthermore, we identify emergent spatio-orientational coupling that is not observed in colonies of straight cells. Our results are important for a fundamental understanding of the interaction and spreading of microorganisms on surfaces, with implications for medical applications and bioengineering.
Figure 1: Illustration of our curved model microorganisms. (a) Each microorganism $i$ is represented by a curved spherocylinder of thickness $d_{0}$ following an arc of radius $R$ and segment angle $\phi_{i}$. We characterize its direction by the unit vector $\hat{\bm{e}}_{i}$ pointing in the direction from the center-of-mass position $\bm{r}_{i}$ radially outward, normal to the arc. The angle $\theta_{i}$ measures the orientation of $\hat{\bm{e}}_{i}$ in the two-dimensional plane. $\bm{r}_{i}^{\mathrm{o}}$ denotes the center-of-arc position for cell $i$. (b) An illustration of calculating the mutual forces of steric interaction between two curved microorganisms. For any two cells $i$ and $j$ in contact, these forces are determined from the overlap of the nearest inner tangent circles of the curved spherocylinders, with center-to-center vector $\bm{d}_{ij}$, using the Hertzian elastic model. (c–d) Illustrative examples of elongated cells for curvatures of (c) $0.011/R=$ 0.01, (d) $0.101/R=$ 0.10, and (e) $0.331/R=$ 0.33. (f) A typical time evolution of the total number of cells $N(t)$ of curvature $0.101/R=$ 0.10 and threshold division length $L_{\mathrm{d}}=5$. The star symbols indicate configurations of $N(t)=2$, $115$, and $1031$, as depicted in panel (g). Color represents the orientation angle $\theta_{i}$ of the cells.
@@
To this end, we introduce a theoretical model to study the growth of the films. It is based on discrete growing mesoscopic curved objects that divide into two halves when their arc length reaches a certain threshold. We illustrate the effects of curvature of the individual microorganisms on the structure of the resulting films. Particular attention is paid to the average size of ordered domains and how it depends on the curvature of the individuals. Moreover, the influence of the threshold length of cell division is investigated. We also discuss the statistical properties of the orientation of curved cells and report a spatio-orientational coupling that we do not observe for straight objects. In addition, we briefly address different structures emerging for growing curved objects.
Materials and Methods
@@
Figure 2: Effects of curvature $1/R$ of the growing individual cells and their threshold division length $L_{\mathrm{d}}$ on the resulting overall structure of the colony. From left to right, curvature increases as $0.011/R=$ 0.01, $0.10$, and $0.33$. From top to bottom, the threshold division length decreases as $L_{\mathrm{d}}=5$, $4$, and $3$. Increasing curvature and decreasing threshold division length lead to finer, less ordered structures and decreasing average domain size. Color represents the angle of orientation $\theta_{i}\in[0,2\pi)$ of each cell $i$, see Fig. 1(a), as quantified by the color bar on the right-hand side. The mean growth rate is set to $g=5\times 10^{-5}$.
@@
Figure 3: (a) Average domain area $\expectationvalue{A}$ as a function of cellular curvature $1/R$ for threshold division lengths $L_{\mathrm{d}}=5$, $4$, and $3$. The average is taken over all domains located within the inner circle of half the colony radius $r_{\mathrm{colony}}$. In the inset we provide a magnification of the regime for large curvature $1/R$ (the dotted rectangle). (b) Dependence of the average domain area for different cellular curvatures $0.011/R=$ 0.01,$0.10$, and $0.33$ on the distance $r$ from the center of the colony. The threshold division length is set to $L_{\mathrm{d}}=5$. We calculate the average $\expectationvalue{A}_{\Delta r}$ for each value of $r$ from all domains located within an annular ring of width $\Delta r=5$. Typically, $r_{\mathrm{colony}}\sim 100$ at times of evaluation. (c) Probability distribution $P(A)$ of the domain area $A$ within colonies of cellular curvature $0.011/R=$ 0.01, $0.10$, and $0.33$. Dashed lines represent exponential behavior $P(A)\sim\exp[-A/A_{*}]$ within the indicated regimes. From corresponding fits, we obtain values of characteristic domain areas $A_{*}=145.4$, $35.2$, and $5.7$, respectively, which decrease with increasing curvature. We set the threshold division length to $L_{\mathrm{d}}=5$. The average is taken over 30 samples.
@@
As outlined above, straight-growing microorganisms are frequently represented by growing spherocylinders that divide into two halves after reaching a certain threshold length [8]. To now consider non-straight microorganisms, we add curvature to these spherocylinders. For this purpose, each curved cell is described as an arc (part of a circle) of radius $R$ and thickness $d_{0}$, see Fig. 1(a). The length of the $i$-th cell is given by the arc length $L_{i}=R\phi_{i}$, where $\phi_{i}$ is the opening angle of the segment of the disk containing the microorganism at its bounding arc. Figures 1(c), (d), and (e) show illustrations of curved cells of curvature $1/R=0.01,0.10,$ and $0.33$.
@@
During the entire process of growth, we need to evaluate the mutual steric interactions resulting from volume exclusion between the cells and their consequences. To this end, at each time step, we find, for any two cells $i$ and $j$ in contact, the two inner tangent circles to the curved spherocylinders that are nearest to each other [14], see Fig. 1(b). The center-to-center vector between these two circles is termed $\bm{d}_{ij}$. We assume the force of mutual interaction $\bm{F}_{ij}$ between the two circles to be described by the Hertzian elastic model [17]. The force $\bm{F}_{ij}$ is then mapped to the interaction between the $i$-th and $j$-th cells.
@@
Here, $\bm{r}_{i}(t)$ denotes the position of the center of mass of the $i$-th cell, and $\theta_{i}(t)$ quantifies the orientation of the $i$-th cell, both at time $t$. The direction of the orientation of the $i$-th cell $\hat{\bm{e}}_{i}=(\cos\theta_{i},\sin\theta_{i})$ is set from the center of the arc to the center-of-mass position of the cell, normal to the arc length, see Fig. 1(a). $\theta_{i}$ takes values in the range $[0,2\pi)$ for curved cells, as curvature provides a sense. $\zeta$ sets the strength of the friction with the substrate per arc length of the cell and is supposed to be identical for all microorganisms. We remark that we assume the friction to be isotropic to concentrate on the steric effects of curvature, while the effect of anisotropic friction was taken into account in previous studies on straight cells [16, 36, 37]. Moreover, $\bm{\ell}_{ij}(t)$ is the vector from the center-of-mass position to the point of the cell where the force acts. $\hat{\bm{e}}_{z}$ is the unit vector pointing into the $z$-direction, normal to the plane.
@@ Results
Figure 4: (a) Schematic illustration of evaluating the spatio-orientational coupling. We determine for each cell $i$ the angle $\gamma_{i}$ for the radial direction running from the center of mass of the colony through the center of mass of the cell $\bm{r}_{i}$. This angle is compared to the individual angular orientation $\theta_{i}$ for each cell. Histograms of $\theta_{i}-\gamma_{i}$ for cellular curvatures (b) $0.011/R=$ 0.01, (c) $0.101/R=$ 0.10, and (d) $0.331/R=$ 0.33 are depicted for the threshold division length $L_{\mathrm{d}}=5$. In (c) and (d), the dashed lines are obtained by fitting the circular normal distribution $\exp[K\cos(\theta-\gamma)]/(2\pi I_{0}(K))$ to the histograms. The parameter $K$ resulting from the fits quantifies the peakedness of the distribution and is depicted in panel (e) as a function of cellular curvature $1/R$ for the different threshold division lengths $L_{\mathrm{d}}=5$, $4$, and $3$.
@@
On this basis, we now explore the fundamental effects of growing and dividing curved microorganisms on the resulting film structures. Videos of the simulated growth process starting from a single, randomly oriented cell under the aforementioned conditions are available in the supporting material. Figure 1(f) shows a typical resulting time evolution of the total number of cells $N(t)$ in a growing colony. It confirms the expected exponential increase. For illustration, Fig. 1(g) provides three intermediate snapshots. Final example configurations upon variation of the basic parameters, that is, curvature $1/R$ [inverse radius, see Fig. 1(a)] and threshold division length $L_{\mathrm{d}}$, are depicted in Fig. 2. All quantities described in the following are analyzed in the final stages of the simulations.
@@
In Fig. 2, color represents the angle of the orientational vector $\theta_{i}\in[0,2\pi)$ of the $i$-th cell, see Fig. 1(a). The columns from left to right show increasing cellular curvature $1/R$. In the left-most column, for the smallest curvature $0.011/R=$ 0.01, the individual cells are basically straight, see Fig. 1(c), and we qualitatively expect corresponding behavior. As already reported in Ref. [8] for straight cells, one does not find global orientational order. Instead, locally aligned domains form a kind of “mosaic” structure. These mesoscopic domains decrease in size as the division length $L_{\mathrm{d}}$ decreases. We recover this trend from top to bottom when we reduce this length.
@@
Our central observation and message is that increasing curvature $1/R$ and decreasing threshold division length $L_{\mathrm{d}}$ generally and substantially reduce the average size of the mesoscopic domains and thus the extent of local alignment. This effect becomes apparent from Fig. 2 when turning from left to right (curvature $1/R$ increasing as $0.01$, $0.10$, $0.33$) and from top to bottom (threshold division length $L_{\mathrm{d}}$ decreasing as $5$, $4$, $3$). Both trends express that straight elongated objects support orientational alignment. Variations by reducing the aspect ratio or deviating from the straight shape lead to less pronounced orientational alignment.
@@
In the following, we analyze these effects more quantitatively. To this end, we determine the average area $\langle A\rangle$ of the mesoscopic domains at any given state of the colony. To determine the sizes of the domains of orientational order, we consider two neighboring cells $i$ and $j$ as belonging to the same mesoscopic domain, if they contact each other and show approximately identical orientation. The first criterion is satisfied if the force in Eq. 4 becomes nonzero. The second criterion applies if their orientation angles fulfill |$\absolutevalue{\theta_{i}-\theta_{j}}<0.1$. (The choice of the precise value of this upper bound for |$\absolutevalue{\theta_{i}-\theta_{j}}$ | does not affect our results qualitatively, see the supporting material). The area $A$ of any resulting mesoscopic domain is calculated as the areal sum over all cells within that domain. We evaluate $\langle A\rangle$ as a function of cellular curvature. Ensemble averages are taken over $10$ different simulations. Figure 3(a) shows the average domain area $\expectationvalue{A}$ as a function of the cellular curvature $1/R$ for the different threshold division lengths $L_{\mathrm{d}}=5,4,$ and $3$. To exclude surface effects, $\expectationvalue{A}$ is calculated in each case for the interior circular film area of half the colony radius $r_{\mathrm{colony}}/2$. Here, the radius of the entire colony is defined via $r_{\mathrm{colony}}(t)=\max_{i=1,...,N}|\bm{r}_{i}(t)|$. The origin is set by the position of the initial cell. In our case, $r_{\mathrm{colony}}\approx 100$, but the precise value slightly depends on the curvature and division length (see the supporting material). From Fig. 3(a), we find that the average domain area $\expectationvalue{A}$ is almost constant when increasing the curvature of the cells $1/R$ at small values, that is, for basically straight microorganisms. Yet, from a certain value of $1/R$ onwards, $\expectationvalue{A}$ significantly decreases. This effect is most pronounced for $L_{\mathrm{d}}=5$, while for $L_{\mathrm{d}}=4$ and $3$ it is more moderate. Since initially the averaged domain size $\langle A\rangle$ is larger for straight cells of $L_{\mathrm{d}}=5$ than for $L_{\mathrm{d}}=4$ and $3$, also the reduction in this value can be larger, as is the case in Fig. 3(a). Still, the magnitudes of $\langle A\rangle$ remain in the order $L_{\mathrm{d}}=5$, $4$, and $3$ from top to bottom. Our observations are in line with the snapshots displayed in Fig. 2. There, we observe by eye reduction in domain size from top to bottom and from left to right in each column and line.
@@
Next, we address in Fig. 3(b) how the area of the domains depends on their location within the colony, together with the effect of cellular curvature on this dependence. We slice the entire colony into concentric, annular rings of width $\Delta r$ for quantification. Their location is given by their distance $r$ from the center of the colony. Figure 3(b) shows the averaged domain area within each ring, $\expectationvalue{A}_{\Delta r}$, as a function of $r$. The ensemble average is taken over 30 different simulations. For basically straight cells of curvature $0.011/R=$ 0.01, we observe a decreasing trend of $\expectationvalue{A}_{\Delta r}$ towards the outer rim of the colony, in line with previous reports in Ref. [8]. Instead, for the increasingly curved cases of $0.101/R=$ 0.10 and $0.33$, $\expectationvalue{A}_{\Delta r}$ remains rather constant towards the outer rim. In line with our results above, we infer that increased curvature alone already implies a significant reduction in domain area and, thus, an increase in mesoscopic disorder. For straight cells, reducing the division length $L_{\mathrm{d}}$ has similar effects [8]. It likewise smoothens the distribution of domain area from the center towards the outer rim of the colony.
@@
Additionally, we evaluate the probability distribution $P(A)$ for the observed mesoscopic domain areas $A$ in a given colony. The distribution shows exponential behavior for straight cells $P(A)\sim\exp[-A/A_{*}]$, where $A_{*}$ represents a characteristic domain area. $A_{*}$ decreases together with the division length $L_{\mathrm{d}}$ [8, 38]. We infer from Fig. 3(c) that the exponential dependence can still be observed with increasing cellular curvature $1/R$. However, increasing curvature reduces the characteristic area $A_{*}$, in line with our qualitative observation in Fig. 2.
@@
Figure 5: Center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ of the microorganisms in the film, see the inset of panel (a), together with color coding of the individual orientations for increasing cellular curvature (a) $0.011/R=$ 0.01, (b) $0.101/R=$ 0.10, and (c) $0.331/R=$ 0.33. The threshold division length is set to $L_{\mathrm{d}}=5$ so that the illustrations reconsider the state of the colonies depicted in the top row of Fig. 2. For the almost straight cells in (a), the radii of the arcs are correspondingly large, therefore several center-of-arc positions $\bm{r}_{i}^{\mathrm{o}}$ are located outside the box. As curvature $1/R$ increases from left to right, the mutual arrangements of $\bm{r}_{i}^{\mathrm{o}}$ transition from (a) rather scattered, gas-like to (c) organized, branched filamentous structures. The latter indicates the stacked organization of curved cells in these colonies.
Thus, as demonstrated above, the curvature of individual cells can have an effect qualitatively similar to reduced division lengths. At least, this concerns the average mesoscopic domain size in the colonies. A similar tendency of reduced domain size could be provoked by reduced aspect ratio of the individual cells by increasing their thickness. Yet, additionally, new structural consequences of cellular curvature emerge that are not present in colonies of straight cells. A specific example is the coupling of cellular orientation to the position within the colony. Through these measures, we can qualitatively and quantitatively distinguish the appearance of colonies of straight cells of low division length from films of curved cells of longer division length, which by eye appear relatively similar; see the bottom left and top right in Fig. 2.
@@
To illustrate this trend, we evaluate their spatio-orientational coupling. We determine for each cell the angle $\gamma_{i}$ of the line running from the center of the colony to the center of mass of the cell. That is, $\gamma_{i}$ measures the radial direction of the colony at the position of the considered cell. We compare this angle with the angular orientation $\theta_{i}$ of each cell; see Fig. 4(a). Proceeding along these lines for all cells, we calculate histograms as a function of $\theta_{i}-\gamma_{i}$. The trend with increasing curvature $0.011/R=$ 0.01, $0.10$, to $0.33$ is illustrated in Fig. 4(b)–(d). The threshold division length here is set to $L_{\mathrm{d}}=5$. Ensemble averages over ten colonies are performed for each curvature.